Vector V⃗ 1 is 6.5 units long and points along the negative x axis. Vector V⃗ 2 is 8.7 units long and points at 60 ∘ to the positive x axis.

a. What are the x and y components of vector V⃗ 1?

b. What are the x and y components of vector V⃗ 2?

c. Determine the magnitude of the sum V⃗ 1+V⃗ 2.

d. Determine the angle of the sum V⃗ 1+V⃗ 2.
Answer #1

Here,

v1 = - 6.5 i

v2 = 8.7 * (cos(60) i + j * sin(60))

v2 = 4.35 i + 7.53 j

a) for the x component and y component of vector V1

x component of v1 = -6.5 unit

y component of v1 = 0 unit

b)

x component of v2 = 8.7 * cos(60)

x component of v2 = 4.35 unit

y component of v2 = 8.7 * sin(60)

y component of v2 =7.53 unit

c)

for the vector v1 + v2

v1 + v2 = -6.5 i + 4.35 i + 7.53 j

v1 + v2 = - 2.15 i + 7.53 j

magnitude of v1 + v2 = sqrt(2.15^2 + 7.53^2)

magnitude of v1 + v2 = 7.84 m/s

the magnitude of v1 + v2 is 7.84 m/s

d)

for the angle of v1 + v2

angle of v1 + v2 = 180 - atan(7.53/2.15)

angle of v1 + v2 = 105.3 degree